Help me finish the vertical analysis schedule

Operating data for Joshua Corporation are presented below. 2020 2019 Sales revenue Cost of goods sold Selling expenses Administrative expenses Income tax expense Net income $748,000 466,752 119,680 64,328 38,148 59,092 $596,000 394,552 75,096 48,872 20,264 57,216 Prepare a schedule showing a vertical analysis for 2020 and 2019. (Round answers to 1 decimal place, eg. 48.5%.)

Amount Percent Amount Percent Sales revenue $748,000 100 $596,000 100 Cost of goods sold 466,752 100 394,552 100 Gross profit 281,248 201,448 Selling expenses 119,680 75,096 Administrative expenses 64,328 48,872 Total operating expenses 184,008 123,968 Income before income taxes 97,240 77,480 Income taxes expense 38,148 20,264 Net income $59,092 $57,216
Answer #1

Percent is calculated =Amount/Sales revenue which is taking sales revenue as base

example

sales revenue%=748000/748000=100%

Cost of Good sold =466752/748000=62.4%

Gross Profit =281248/748000=37.6%

FYrGCUf8PmYeFKWnRAgkAAAAASUVORK5CYII=